OpenStudy (anonymous):

A car has an engine which delivers a constant power. It accelerates from rest at time t = 0, and at t = t0 its acceleration is a0. What is its acceleration at t = 2t0? Ignore energy loss due to friction.

OpenStudy (anonymous):

constant power meaning, constant force acting .. so i guess acceleration remains unchanged?!

OpenStudy (anonymous):

.nope. got the answer :) P = FV = ma*at = ma^2 t . therefore m(a0)^2t = m(x)^2 (2t) solve for x which is the acceleration at 2t. answer is a0/sqrt(2). mission achieved :) intuitively, acceleration needs to be negative, else the engine would go one forever.

OpenStudy (anonymous):

go on accelerating forever that is. it needs to reach a "carrying capacity"

OpenStudy (anonymous):

 so u are right